# On Emergent Social World Models — Evidence for Functional Integration of Theory of Mind and Pragmatic Reasoning in Language Models

### v1

@@ 1 Introduction
Large language models (LMs) possess astonishing abilities and prove useful for a plethora of downstream tasks, but controversy persists regarding how to conceptualize their capacities. Some view LMs as mere “stochastic parrots,” fundamentally incapable of genuine understanding Bender and Koller (2020); Bender et al. (2021). Others argue that highly proficient language ability presupposes the emergence of “world models,” i.e., representations of (relevant aspects) of the language-generating processes Yudkowsky (2023); Bereska and Gavves (2024). Arguments for latter view typically come from models optimized for solving finite-state problems like Chess Toshniwal et al. (2022), Othello Li et al. (2023), or taxi navigation Vafa et al. (2024), or from narrow domains like text-based adventure games Li et al. (2021). This paper contributes to the debate by investigating the extent to which large pretrained LMs recruit representations and computational mechanisms suggestive of “world models” for tasks crucial for general-purpose language understanding.
@@
Language understanding comprises formal and functional abilities Mahowald et al. (2024). Formal abilities (e.g., syntax) may be acquired relatively straightforwardly from local statistical patterns in training data. Functional abilities relate to language use, e.g., pragmatic reasoning, such as comprehension of contextual meaning and social implications of utterances Levinson (1983). Functional ability seems more difficult to extract from text alone, since the surface language signal provides less direct information about relevant latent variables, such as interlocutors’ mental states. This raises the issue: to what extent LMs develop rich “social world models”—entertaining concepts of self and other, social commitment, and nested beliefs—in order to achieve their observed level of performance? We here address this big-picture question with a concrete case study on the relationship between LMs’ capacities for (i) Theory of Mind reasoning and (ii) pragmatic reasoning.
@@
Theory of Mind (ToM) refers to the capability to attribute mental states—unobservable, latent properties such as beliefs or desires—to oneself and other agents, and plays a fundamental role in social cognition, enabling individuals to interpret others’ behavior as causally driven by these mental states (Wimmer and Perner, 1983; Beaudoin et al., 2020). A longstanding tradition in philosophy, linguistics, and cognitive science conceptualizes linguistic behavior as a specialized form of goal-directed action Grice (1975), such that understanding language is, in principle, structurally similar to interpreting non-linguistic actions. Indeed, evidence from cognitive neuroscience indicates that overlapping neural substrates support both Theory of Mind and pragmatic reasoning Saxe and Kanwisher (2003); Bosco et al. (2017); Hauptman et al. (2023).
@@
What holds for human cognition need not, of course, generalize to LMs. Nothing in their architecture explicitly enforces structural similarity to human solutions for functional language abilities. Nevertheless, if shared computational principles govern linguistic and non-linguistic social reasoning in humans, efficient compression principles might favor cross-purposing internal mechanisms across these domains in stochastic learners as well. We therefore formulate two contrasting hypotheses. The functional specialization hypothesis holds that LMs use different computational mechanisms for pragmatic and ToM-reasoning. In contrast, the functional integration hypothesis maintains that computational mechanisms for pragmatic and ToM-reasoning are shared, at least in substantial part.
@@
This dichotomy structures our investigation, even if itis a simplification in several respects. The repurposing of computational mechanisms likely exists along a continuum rather than as a binary distinction, and the constructed domains of pragmatic and ToM-reasoning may themselves lack internal homogeneity. To acknowledge this general question of the level of granularity at which function sharing should be investigated, we build on recent like-minded work Ma et al. (2023) and investigate LMs’ ToM capabilities while taking into account the ATOMS framework, short for “Aspects in Theory Of Mind Space,” which identifies seven subcategories in ToM abilities (Beaudoin et al., 2020) (described in more detail in Appendix A.1).
@@
We explore evidence from behavioral evaluations and causal-mechanistic experiments (Sections 2 & 3 respectively). The latter builds on recent work by AlKhamissi et al. (2025a) on functional localization inspired from cognitive neuroscience (cf., Duan et al., 2025; Hanna et al., 2025). While this work did not find strong evidence for functional ToM localization based on a very small data set, we do find causal evidence for it based on a much larger data set which is part of this paper’s contribution, as is a more thorough investigation of the method of functional localization in LMs itself. Our paper contributes to related work separately focusing on ToM and linguistic abilities, in the domain of mechanistic interpretability of LMs (Wang et al., 2022), in particular their language capabilities (Oh et al., 2025) and mental state representations (Bortoletto et al., 2024; Zhu et al., 2024), and to the growing number of behavioral evaluations of LMs’ ToM and pragmatic abilities (which, in part, provided mixed results; e.g., Hu et al., 2023; Ullman, 2023; Sap et al., 2022; Ma et al., 2025). While the functional connection of these two abilities has been addressed more in human research (Baldwin and Saylor, 2005; Hauptman et al., 2023), work on the functional organization of LMs has mostly focused on aligning only LMs’ language representations to human brains (Tuckute et al., 2024; AlKhamissi et al., 2025b). We follow recent arguments for more integrated investigations of linguistic abilities of LMs grounded in cognitive science approaches (Sumers et al., 2023; McCoy et al., 2024). In sum, this paper contributes behavioral and causal-mechanistic experiments, paired with stringent hypothesis-driven statistical testing, that provide suggestive evidence in favor of the functional integration view, thus contributing to the question of emergent social world models. In doing so, the paper contributes novel, empirically grounded ToM localizer data and methodological refinements of the functional localization method. [^1] We release all code and data at [redacted for reviewing].
@@ 2.1 Datasets
We select 16 datasets testing ten different pragmatic phenomena from extant pragmatics benchmarks (Ma et al., 2025), and 22 datasets from existing ToM benchmarks. We expand upon fine-grained annotations of ToM benchmarks introduced by Ma et al. (2023) for more ToM datasets, who annotate LM benchmarks with ATOMS identified by Beaudoin et al. (2020) that the benchmarks likely test. The detailed list of datasets and annotations can be found in Appendix A. All datasets have a multiple-choice format: for each item $x$, given a context $c$ that contains the target vignette, a question, and optionally instructions or the answer options, there are $2\leq n\leq 6$ answer options, one of which is the correct answer that is consistent with ToM-based or pragmatic reasoning. The full prompts for each dataset are reported in Appendix A.3. For evaluating general language ability the BLiMP dataset is used (Warstadt et al., 2020).
@@ 2.2 Models & Procedure
We test 48 open-source LMs from seven families (Llama, Qwen, Falcon, Mistral, OLMo, Pythia, Gemma), ranging from 0.5B to 72B parameters (Appendix A.4). Where possible, we test base and fine-tuned versions, the type of which (instruction-tuning, RLHF, both) depends on the model.
@@ 2.3 Results
To address prediction P2 and explore models’ performance more closely, we fit a Bayesian beta regression model in R using the brms package (Bürkner, 2017), regressing the average accuracy of each model on each dataset against the main effects of model family (seven levels), model size (where models up to 8B parameters were labeled as small, up to 32B as medium, and up to 72B as large), model type (base vs. fine-tuned), dataset type (whether instruction or answer options were in context) and domain (ToM vs. pragmatics), an interaction between model size and type, and random by-dataset intercepts. [^2] The regression model in R syntax: accuracy $\sim$ model_family + size * model_type + ds_type + domain + (1 | dataset). A model with full interaction structure did not converge. Here and below, we report posterior means and 95% credible intervals for the posterior of contrasts of interest. All independent variables were sum-coded. In line with most evaluations in the literature, we found that, across domains, fine-tuned models performed credibly better than base models ($\beta=0.11[0.05,0.16]$), large models performed better than medium models ($\beta=0.20\;[0.11,0.29]$) and medium models performed better than small models ($\beta=0.26[0.20,0.32]$). We found no differences between the dataset types ($\beta=-0.06[-0.88,0.73]$). Crucially, in line with prediction P2, we also found no difference between the domains ToM and pragmatics ($\beta=-0.03[-0.74,0.67]$), suggesting that the information about the domain of a given task is not crucial for predicting a given model’s accuracy if sufficient information about the model is given. Further accuracy results on the single datasets are in the Appendix (Figure 5 and Figure 6).
@@
Finally, we address P3 with model comparison for two Bayesian beta regression models predicting the average accuracy on pragmatics datasets for each model against the baseline predictors of model size and model type. [^3] Including the model family fixed effect led to convergence issues. One model (M0) contains the additional predictor of the model’s average accuracy on the syntax benchmark. The other model (M1) contains information of the model’s accuracy on ToM tasks. Comparing M0 and M1 via leave-one-out cross-validation (LOO-CV), we find that model M1 was significantly better than M0 ($ELPD=-16.8,\;p\text{-value}=0$), thus supporting prediction P3.
@@
Conjointly, these behavioral results provide modest indirect evidence for the functional integration hypothesis. But to explore potential diversity in ToM reasoning further, we also ask whether there are more specific aspects of the ToM capabilities that LMs perform particularly well on. We operationalize this question through our annotations of the evaluation datasets with the ATOMS framework Beaudoin et al. (2020). We compare the baseline beta regression model (MB) where a model’s average accuracy on ToM evaluation datasets is regressed against the model family, size, type and dataset type, to models which additionally include all possible combinations of the seven ATOMS main effects (belief, desire, intention, knowledge, non-literal communication, emotion, percepts), resulting in 128 different regression models. Comparison was also conducted via LOO-CV. We found that the baseline MB was numerically worse than any model with ATOMS predictors. The best model additionally contained the aspects intentions, desires, emotions, percepts and non-literal communication. The difference between MB and the best model was significant ($ELPD=-58.80\;p\text{-value}=0$; full results are in the Appendix, Figure 7). Model comparison further suggests that several models are on par with the best model, and several models next to MB are clearly worse than others. The least performing models do not contain the predictor percepts, while containing up to three additional other ATOMS predictors. This suggests that the factor percepts added the most predictive power to the statistical models of the LMs’ accuracy, specifically distinguishing datasets which tapped into this ToM aspect (datasets containing false-belief tasks about the location of an object and tasks about agent properties) from other evaluation datasets.
@@ 3 Functional Localization
Functional subnetwork localization is a method that has been widely used in cognitive neuroscience to identify regions in the human brain that are implied in the execution of a particular task or ability, including Theory of Mind (e.g., Saxe et al., 2006; Fedorenko et al., 2010), and indeed Hauptman et al. (2023) suggests that pragmatic language use in humans relies on the same ToM subnetwork. The procedure uses so-called localizer suites that contrast a target condition which contains stimuli that require the capacity of interest for processing, against a closely matched control condition which does not require it. For example, the ToM network has been localized by contrasting stories in which an actor holds a false belief (Baron-Cohen et al., 1985) against stories about pictures containing false information (Zaitchik, 1990). While reading these stories, the brain activity of human participants is recorded and localization proceeds via by-individual identification of regions that are selectively active under the target, but not under the control stimuli. For our localizations, we build on recent work extending this procedure to language models (AlKhamissi et al., 2025a). However, given that ToM may itself not be a single monolithic capability (see Section 1), cognitive neuroscience has also used other localizers than false belief tasks, e.g., economic games or stories involving moral intent, which have been found to activate the same functional brain region in humans.
LatentBeliefs (Saxe and Kanwisher, 2003)
@@
GameBeliefs (Chang et al., 2023)
@@
Table 1: Example synthetic stimuli for five conditions from two localizer suites. Dots indicate answer options. “S / Q:” stands for “Statement / question:” used in the prompts. Examples for all suites are in the appendix, Table 3.
@@ 3.1 Localizer Suites
A localizer suite $L$ comprises a set of sets of target stimuli $\{S_{T}^{(j)}\}_{1\leq j\leq n_{j}}$ and a set of sets of control stimuli $\{S_{C}^{(k)}\}_{1\leq k\leq n_{k}}$. Localizer suites with $n_{j}=n_{k}=1$ are called simple. We created four localizer suites based on extant experimental material which was successful in previous fMRI studies for singling out the ToM network in human brains. Prior studies were selected for variety and to cover all of the seven top-level categories of the ATOMS framework, while still sticking with empirically validated material: [^4] For example, belief and percepts are relevant for the LatentBelief suite, intentions and knowledge are relevant for the MoralIntent suite, desires and emotions are relevant for the GameBelief suite, and the CommunicativeIntent suite addresses non-literal communication. The LatentBeliefs suite covers the commonly used false-belief items Saxe and Kanwisher (2003), CommunicativeIntent looks at mental state ascriptions in communication, and so might be particularly relevant for comparison to pragmatic reasoning Bosco et al. (2017), GameBeliefs considers ToM reasoning in the context of economic games (trust & ultimatum; Chang et al., 2023), and MoralIntent considers issues of responsibility in decision making Young et al. (2007). Localizer suites contained different numbers of target and control sets (see Table 1 for overview and examples). The two simple suites (GameBeliefs and MoralIntent) have target and control sets in which items are exactly paired. To mitigate data contamination and to control for irrelevant surface-level variation in the localizer materials that may affect LMs more than humans (e.g., differences in semantic content of target vs. control stimuli), we used GPT-5 (which is not evaluated in this study) with a 5-shot in-context prompt to generate 100 novel synthetic stimuli for each dataset in each condition (see Appendix B for prompts used). All our items ended in a task question, which would require a human participant to engage in Theory of Mind reasoning in the target stimuli. More information about the data is given in Appendix C, with details on prompt formatting in Section C.2.

### v2

@@ 1 Introduction
Large language models (LMs) possess astonishing abilities and prove useful for a plethora of downstream tasks, but controversy persists regarding how to conceptualize their capacities. Some view LMs as mere “stochastic parrots,” fundamentally incapable of genuine understanding Bender and Koller (2020); Bender et al. (2021). Others argue that highly proficient language ability presupposes the emergence of “world models,” i.e., representations of (relevant aspects of) the language-generating processes Yudkowsky (2023); Bereska and Gavves (2024). Arguments for the latter view typically come from models optimized for solving finite-state problems like Chess Toshniwal et al. (2022), Othello Li et al. (2023), or taxi navigation Vafa et al. (2024), or from narrow domains like text-based adventure games Li et al. (2021). But to what extent might large pretrained LMs recruit representations and computational mechanisms suggestive of “world models” for tasks crucial for general-purpose language understanding?
@@
Language understanding comprises formal and functional abilities Mahowald et al. (2024). Formal abilities (e.g., syntax) may be acquired relatively straightforwardly from local statistical patterns in training data. Functional abilities relate to language use in communication, e.g., pragmatic reasoning, such as comprehension of contextual meaning and social implications of utterances Levinson (1983). Functional ability seems more difficult to extract from text alone, since the surface language signal provides less direct information about relevant latent variables, such as interlocutors’ mental states. This raises the issue: to what extent do LMs develop rich “social world models”—entertaining concepts of self and other, social commitment, and nested beliefs—in order to achieve their observed level of performance?
@@
We here address this big-picture question with a concrete case study on the relationship between LMs’ capacities for (i) Theory of Mind reasoning and (ii) pragmatic reasoning. Theory of Mind (ToM) refers to the capability to attribute mental states—unobservable, latent properties such as beliefs or desires—to oneself and other agents, and plays a fundamental role in social cognition (Beaudoin et al., 2020) and pragmatic language use (Grice, 1975; Goodman and Frank, 2016). Evidence from cognitive neuroscience indicates that humans indeed have, at least to some extent, shared functional mechanisms, namely, overlapping neural substrates supporting both Theory of Mind and pragmatic reasoning Saxe and Kanwisher (2003); Bosco et al. (2017); Hauptman et al. (2023).
@@
What holds for human cognition need not, of course, generalize to LMs. Nothing in their architecture explicitly enforces structural similarity to human solutions for functional language abilities. Nevertheless, if shared computational mechanisms govern linguistic and non-linguistic social reasoning in humans, efficient compression principles might favor cross-purposing internal mechanisms across these domains in stochastic learners as well. We therefore formulate two contrasting hypotheses. The functional specialization hypothesis holds that LMs use different computational mechanisms for pragmatic and ToM-reasoning. In contrast, the functional integration hypothesis maintains that computational mechanisms for pragmatic and ToM-reasoning are shared, at least in substantial part.
We explore whether LMs exhibit functional specialization or functional integration through evidence from behavioral evaluations and causal-mechanistic experiments. The latter builds on recent work by AlKhamissi et al. (2025a) on functional localization inspired from cognitive neuroscience which identifies subsets of neurons in LMs supporting performance on target tasks (cf., Duan et al., 2025; Hanna et al., 2025). While their work did not find strong evidence for functional ToM localization based on a very small data set, we do find causal evidence for it based on a much larger data set which is part of this paper’s contribution.
@@
The two hypotheses structure our investigation, even if this dichotomy is a simplification in several respects. The repurposing of computational mechanisms likely exists along a continuum rather than as a binary distinction, and the constructed domains of pragmatic and ToM-reasoning may themselves lack internal homogeneity. To acknowledge this general question of the level of granularity at which function sharing should be investigated, we investigate LMs’ ToM capabilities while taking into account the “Aspects in Theory Of Mind Space” (ATOMS) framework. ATOMS identifies seven fine-grained subcategories in ToM abilities (Beaudoin et al., 2020; beliefs, intentions, desires, emotions, knowledge, percepts, mentalistic understanding of non-literal communication; described in detail in Appendix A.1). We conduct evaluations of LMs on datasets tapping into different ATOMS, building on like-minded work Ma et al. (2023).
Related Work & Contributions
Our paper builds on ideas from numerous studies in psychology and cognitive science, showing that Theory of Mind enables humans to interpret others’ actions as causally driven by these mental states (e.g., Beaudoin et al., 2020; Goodman et al., 2006; Wimmer and Perner, 1983). A longstanding tradition in philosophy, linguistics, and cognitive science conceptualizes communication as a specialized form of goal-directed action Austin (1975); Grice (1975), such that understanding language is, in principle, structurally similar to interpreting non-linguistic actions, and has also been conceptualized as interlocutors engaging in probabilistic recursive reasoning about each other’s mental states (Grice, 1975; Goodman and Frank, 2016).
@@
Additionally, we contribute to related work evaluating LMs, often separately focusing on ToM and linguistic abilities, in the domain of mechanistic interpretability (Wang et al., 2022), in particular LMs’ language capabilities (Oh et al., 2025) and mental state representations (Bortoletto et al., 2024; Zhu et al., 2024), and to the growing number of behavioral evaluations of LMs’ ToM and pragmatic abilities (which, in part, provided mixed results; e.g., Hu et al., 2023; Ullman, 2023; Sap et al., 2022; Ma et al., 2025). While the functional connection of these two abilities has been addressed more in human research (Baldwin and Saylor, 2005; Hauptman et al., 2023), work on the functional organization of LMs has mostly focused on aligning only LMs’ language representations to human brains (Tuckute et al., 2024; AlKhamissi et al., 2025b). We follow recent arguments for more integrated investigations of linguistic abilities of LMs grounded in cognitive science approaches (Sumers et al., 2023; McCoy et al., 2024).
In sum, this paper contributes behavioral and causal-mechanistic experiments (Sections 2 & 3 respectively), paired with stringent hypothesis-driven statistical testing, that provide suggestive evidence in favor of the functional integration view, thus contributing to the question of emergent social world models. In doing so, the paper contributes novel, empirically grounded ToM localizer data and methodological refinements of the functional localization method. [^1] We release all code and data at https://github.com/polina-tsvilodub/lm-emergent-social-world-models.
@@ 2 Behavioral Evaluation
While P1–P2 address the behavioral correlation between ToM and pragmatic capabilities, P3 helps to disentangle this connection from potential confounds: both ToM and pragmatic capabilities likely correlate with general language ability (e.g., for instruction comprehension), as well as on general cognitive abilities and world knowledge. While the following analyses focus on controlling for general linguistic ability through performance on two linguistic benchmarks, we provide additional exploratory mechanistic analyses of the models also on general intelligence tasks like analogical reasoning and entity tracking in Section 3.5 and in Appendix D.1.
@@ 2.1 Datasets
We select 16 datasets testing ten different pragmatic phenomena from extant pragmatics benchmarks (Ma et al., 2025), and 22 datasets from existing ToM benchmarks. We use several ToM benchmarks with ATOMS annotations introduced by Ma et al. (2023), and select additional ToM datasets, annotating them with ATOMS we consider the benchmarks to likely test. All datasets have a multiple-choice format: for each item $x$, given a context $c$ that contains the target vignette, a question, and optionally instructions or the answer options, there are $2\leq n\leq 6$ answer options, one of which is the correct answer that is consistent with ToM-based or pragmatic reasoning. For evaluating general language ability of each model, the BLiMP dataset (Warstadt et al., 2020) and the SNLI dataset (Bowman et al., 2015) are used. The detailed list of datasets and annotations can be found in Appendix A, full prompts for each dataset are reported in Appendix A.3.
@@ 2.2 Models & Procedure
We test 48 open-weights and/or open-source LMs from seven families (Llama, Qwen, Falcon, Mistral, OLMo, Pythia, Gemma), ranging from 0.5B to 72B parameters (listed in Appendix A.4). Where possible, we test base and fine-tuned versions, the type of which (instruction-tuning, RLHF, both) depends on the model.
@@ 2.3 Results
To address prediction P2, we fit a Bayesian beta regression model in R using the brms package (Bürkner, 2017), regressing the average accuracy of each model on each dataset against the main effects of model family (seven levels), model size (where models up to 8B parameters were labeled as small, up to 32B as medium, and up to 72B as large), model type (base vs. fine-tuned), dataset type (instruction / answer options in context vs. sentence completion) and domain (ToM vs. pragmatics), an interaction between model size and type, and random by-dataset intercepts. [^2] The regression model in R syntax: accuracy $\sim$ model_family + size * model_type + ds_type + domain + (1 | dataset). A model with full interaction structure did not converge. Here and below, we report posterior means and 95% credible intervals for the posterior of contrasts of interest. All independent variables were sum-coded. In line with most evaluations in the literature, we found that, across domains, fine-tuned models performed credibly better than base models ($\beta=0.11[0.05,0.16]$), large models performed better than medium models ($\beta=0.20[0.11,0.29]$) and medium models performed better than small models ($\beta=0.26[0.20,0.32]$). We found no differences between the dataset types ($\beta=-0.06[-0.88,0.73]$). Crucially, in line with prediction P2, we also found no difference between the domains ToM and pragmatics ($\beta=-0.03[-0.74,0.67]$), suggesting that the information about the domain of a given task is not crucial for predicting a given model’s accuracy if sufficient information about the model is given. Further accuracy results on the single datasets are in the Appendix (Figure 5–Figure 8).
@@
Finally, we address P3 with model comparison for two Bayesian beta regression models predicting the average accuracy on pragmatics datasets for each model against the baseline predictors of model size and model type. [^3] Including the model family fixed effect led to convergence issues. One model (M0) contains the additional predictor of the model’s average accuracy on the linguistic benchmarks. The other model (M1) contains information of the model’s average accuracy on ToM tasks. Comparing M0 and M1 via leave-one-out cross-validation (LOO-CV), we found that model M1 was significantly better than M0, measured through expected log predictive density (ELPD) ($ELPD=-16.1,\;p\text{-value}=0$), thus supporting prediction P3. Additional details on our statistical analyses are in Appendix A.5.
@@
To explore potential diversity in ToM reasoning further, we also ask whether there are more specific aspects of the ToM capabilities that LMs perform particularly well on. We operationalize this question through our annotations of the evaluation datasets with the ATOMS framework Beaudoin et al. (2020). We compare the baseline beta regression model (MB) where a model’s average accuracy on ToM evaluation datasets is regressed against the model family, size, type and dataset type, to models which additionally include all possible combinations of the seven ATOMS main effects, resulting in 128 different regression models. Comparison was also conducted via LOO-CV. We found that the baseline MB was numerically worse than any model with ATOMS predictors. The best model additionally contained the aspects intentions, desires, emotions, percepts and non-literal communication. The difference between MB and the best model was significant ($ELPD=-58.80\;p\text{-value}=0$; full results are in the Appendix, Figure 9). Model comparison further suggests that several models are on par with the best model, and several models next to MB are clearly worse than others. The least performing models do not contain the predictor percepts, while containing up to three additional other ATOMS predictors. This suggests that the factor percepts added the most predictive power to the statistical models of the LMs’ accuracy, specifically distinguishing datasets which tapped into this ToM aspect (false-belief tasks about the location of an object and tasks about agent properties) from other evaluation datasets. Conjointly, these behavioral results provide first evidence for the functional integration hypothesis. Next, we investigate whether the hypothesis is supported at the level of LMs’ internal mechanisms.
@@ 3 Functional Localization
Functional subnetwork localization is a method that has been widely used in cognitive neuroscience to identify regions (or, subnetworks) in the human brain that are implied in the execution of a particular task or ability, including Theory of Mind (e.g., Saxe et al., 2006; Fedorenko et al., 2010), and indeed Hauptman et al. (2023) suggest that pragmatic language use in humans relies on the same ToM subnetwork. The procedure uses so-called localizer suites that contrast a target condition which contains stimuli that require the capacity of interest for processing, against a closely matched control condition which does not require it. For example, the ToM network has been localized by contrasting target stories in which an actor holds a false belief (Baron-Cohen et al., 1985) against control stories about pictures containing false information (Zaitchik, 1990). While reading these stories, the brain activity of human participants is recorded and localization proceeds via by-individual identification of regions that are selectively active under the target, but not under the control stimuli. However, given that ToM may itself not be a single monolithic capability (see Section 1), cognitive neuroscience has also used other localizers than false belief tasks, e.g., economic games or stories involving moral intent, which have been found to activate the same functional brain region in humans. For our localizations, we build on recent work extending this procedure to language models (AlKhamissi et al., 2025a), described next.
LatentBeliefs (Example of GPT-5 generated stimuli based on data from Saxe and Kanwisher, 2003)
@@
GameBeliefs (Example of GPT-5 generated stimuli based on data from Chang et al., 2023)
@@
Table 1: Example synthetic stimuli for five conditions from two localizer suites. Dots indicate answer options. “S” / “Q:” stand for “Statement” / “question:” used in the prompts. “trgt” indicates a target stimulus, “ctrl” a control stimulus. Examples for all suites are in the Appendix, Table 3.
@@ 3.1 Localizer Suites
A localizer suite $L$ comprises a set of sets of target stimuli $\{S_{T}^{(j)}\}_{1\leq j\leq n_{j}}$ and a set of sets of control stimuli $\{S_{C}^{(k)}\}_{1\leq k\leq n_{k}}$. Localizer suites with $n_{j}=n_{k}=1$ are called simple. We synthetically generated four localizer suites based on extant experimental material which was successful in previous fMRI studies for singling out the ToM network in human brains. These prior studies were selected for variety and to cover all of the seven top-level categories of the ATOMS framework, while still sticking with empirically validated material: [^4] For example, belief and percepts are relevant for the LatentBelief suite, intentions and knowledge are relevant for the MoralIntent suite, desires and emotions are relevant for the GameBelief suite, and the CommunicativeIntent suite addresses non-literal communication. The LatentBeliefs suite covers the commonly used false-belief items Saxe and Kanwisher (2003), CommunicativeIntent looks at mental state ascriptions in communication, and might be particularly relevant for comparison to pragmatic reasoning Bosco et al. (2017), GameBeliefs considers ToM reasoning in the context of economic games (trust & ultimatum; Chang et al., 2023), and MoralIntent considers issues of responsibility in decision making Young et al. (2007). Our localizer suites contain different numbers of target and control sets (each called condition): the two simple suites (GameBeliefs and MoralIntent) have target and control sets in which items are exactly paired, two suites (LatentBeliefs and CommunicativeIntent) have several unpaired sets (see Table 1 for examples).
To mitigate data contamination and to control for irrelevant surface-level variation in the localizer materials that may affect LMs more than humans (e.g., differences in semantic content of target vs. control stimuli), we used GPT-5 (not evaluated in this study) with a 5-shot in-context prompt to generate 100 novel synthetic stimuli for each condition in each suite (see Appendix B for all prompts). To critically assess the quality of the novel synthetic stimuli against the established original stimuli, we ran a Principal Component Analysis with two components comparing sentence embeddings from SentenceTransformer’s all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019) of the original and synthetic stimulus stories. For all conditions except HumanDescr from the LatentBeliefs suite, we confirmed that original and synthetic stimuli did not form clearly separate clusters, indicating that the synthetic stimuli are overall comparable to the original ones. For the more divergent HumanDescr condition, the most likely reason is that the original stimuli only described the physical appearance of humans, while the synthetic stimuli covered other aspects as well. This difference, however, does not impact their role as a non-ToM control condition. More details and visualizations of the PCA results are in Appendix C.3.
All localizer items consist of context sentences and end in a task question, which would require a human participant to engage in Theory of Mind reasoning in the target conditions, and multiple choice answer options. More details about all localizer suites are given in Appendix C.